The Coffee Market Price Per Pound Quantity Demanded Quantity Supplied $2.00 25 0 $3.00 20 3...

  1. The Coffee Market

Price Per Pound

Quantity Demanded

Quantity Supplied

$2.00

25

0

$3.00

20

3

$4.00

12

5

$5.00

10

10

$6.00

6

15

$7.00

3

20

  1. Graph the supply and demand curves above on a market model.
  2. What would be the equilibrium price in the market? Show this on your market model.
  3. If the price was set at $3.00, what would you observe in the coffee market? Show this situation on your market model.
Answer #1

Answer

a)

To graph the model we will place the points on the curves and join it as follows.

B)

The equilibrium is at Qd=Qs=10 units where P=$5

C)

P=$3

Qd=20 and Qs=3 units

Qd>Qs so there is a shortage

Shortage =Qd-Qs=20-3=17 units